OpenStudy (anonymous):

Tims average for the past five tests is 84. To increase his average to 86, what must he make on the sixth test?

OpenStudy (anonymous):

oops 96

OpenStudy (anonymous):

5x84 + x = 86*6

OpenStudy (anonymous):

84 * 5 = 340 (340 + x) / 6 = 86 340 + x = 516 x = 176

OpenStudy (anonymous):

:x that loooks wrong... LOL

OpenStudy (anonymous):

Wow. Thanks for all the answers guys. Let me read em' and I'll give you medals. ;D

OpenStudy (anonymous):

x = 96

OpenStudy (anonymous):

84 x 5 = 420 :)

OpenStudy (anonymous):

Okay. I'm confused. Everyone has different answers... ._.

OpenStudy (anonymous):

ahhh thanks LOL 84 * 5 = 420 (420 + x) / 6 = 86 420+ x = 516 x = 96

OpenStudy (anonymous):

its 96

OpenStudy (anonymous):

it's 96 sorry

OpenStudy (anonymous):

lol. 96. K. Thanks guys.

OpenStudy (lila):

A sample of legal costs ( in thousands of dollars) for school districts for two recent consecutive years is shown. At a = 0.05 is there a difference in the costs? Suggest a reason for the difference, if one exists. Year 1 / 108, 36,65, 108, 87, 94, 10, 40 Year 2 / 138, 28, 67, 181, 97, 126, 18, 67 please solve